OpenStudy (melstutes):

Which of the following equations is the equation of a suitable line of best fit for the data given below? {(45, 18), (62, 13), (72, 11), (55, 15), (38, 21), (26, 31), (93, 9), (64, 13), (29, 28), (33, 25), (49, 17)} A. y = 3x + 35 B. y = –10x + 35 C. - 3/10x -35 D. -3/10x + 35

OpenStudy (melstutes):

PLEASE help

OpenStudy (anonymous):

I'd go with D, although I'm not sure... The pattern works with D's equation..It is precise but not completely accurate (y values are off a little)

OpenStudy (kropot72):

Should the options C and D be written: y = -3/10x - 35 y = - 3/10x + 35

OpenStudy (anonymous):

@melstutes I based my answer on C and D being as @kropot72 said. Are they supposed to be like that?